In 2015, jim earnss $100, milk costs $2 per litre, and cookies cost $4 per dozen. Draw Jim's budget constraint.



1
Expert's answer
2022-02-16T13:08:22-0500

"PA\times QA + PB \times QB = M"


where PA is price of milk and PB is price of cookies


"milk \times 2 dollars + cookies \times 4 dollars = 100 dollars"

let quantity of milk be y and cookies be x

"100=4x+2y"

"y=-2x+50"